OpenStudy (anonymous):

A confidence interval for a population mean has a margin of error of 3.2 ....determine the length of the interval

jimthompson5910 (jim_thompson5910):

length of interval = 2*(margin of error) length of interval = 2*3.2 length of interval = 6.4